OpenStudy (anonymous):

Use the equation and type the ordered-pairs. y = log 3 x {(1/3,__), (1,__), (3,__), (9,__), (27,__), (81,__)}

OpenStudy (anonymous):

\[y =\log_{3} x\]

OpenStudy (shamim):

I wanna show u last 1

OpenStudy (shamim):

U know x=81, y=?

OpenStudy (anonymous):

okay..

OpenStudy (shamim):

y=log3(81)=log3(3^4)=?

OpenStudy (anonymous):

where did you get the "4"?

OpenStudy (shamim):

Ok tell me 3^2=?

OpenStudy (anonymous):

9

OpenStudy (shamim):

3^3=?

OpenStudy (shamim):

3^4=?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

81

OpenStudy (shamim):

So we can write 81=3^4 Right?!

OpenStudy (anonymous):

right.

OpenStudy (shamim):

So we can write log3(81)=log3(3^4) ?

OpenStudy (anonymous):

yep

OpenStudy (shamim):

Can u tell me logx(x^2)=?

OpenStudy (anonymous):

not really. no.

OpenStudy (shamim):

I think u know logx(x)=1 right?

OpenStudy (shamim):

Log2(2)=1 Right?

OpenStudy (shamim):

Lig3(3)=1 Right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

log2(2) = 0.602059991328

OpenStudy (shamim):

This is law of logarithm

OpenStudy (shamim):

Log4(4)=1

OpenStudy (shamim):

Logy(y)=1 Right?!

OpenStudy (anonymous):

i guess. although my calculator doesnt agree.

OpenStudy (shamim):

Ofcourse ur calculator will b agree

OpenStudy (shamim):

Test log3(3)=?

OpenStudy (shamim):

Anyway it is 3 basis log 3

OpenStudy (shamim):

Plz tell me the result frm ur calculator

OpenStudy (anonymous):

1.43136376416

OpenStudy (shamim):

Never

OpenStudy (shamim):

U r doing something wrong!

OpenStudy (shamim):

log3(3)=1

OpenStudy (anonymous):

yep, i found out i was putting (log(3)) * 3

OpenStudy (anonymous):

anyways, continue.

OpenStudy (shamim):

Law of logarithm is log3(3^2)=2log3(3)=?

OpenStudy (anonymous):

2

OpenStudy (shamim):

Correct!!!!

OpenStudy (shamim):

log3(3^3)=3*log3(3)=?

OpenStudy (shamim):

Tell me log3(81)=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

4

OpenStudy (shamim):

This is ur result!!!

OpenStudy (anonymous):

then i am guessing that these are all my answers... 1/2; 0; 1; 2; 3; 4?

OpenStudy (shamim):

81,4

OpenStudy (anonymous):

{(1/3,1/2), (1,0), (3,1), (9,2), (27,3), (81,4)}?

OpenStudy (anonymous):

i may be wrong about the 9

OpenStudy (shamim):

No 9,2 is correct!

OpenStudy (anonymous):

so the 1/2 is wrong... is should be -1?

OpenStudy (shamim):

Ya

OpenStudy (shamim):

1/3,-1 is correct!

OpenStudy (anonymous):

i guess that is all i needed then.

OpenStudy (shamim):

Ok. Thank u !

OpenStudy (anonymous):

and thank you.